• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name averett

    Log entry time 00:00:49 on August29,2003

    Entry number 114113

    keyword=swing shift summary



    16:30 Data from recent runs indicate target position is incorrect. Also MCC has trouble clearing FSD with target in position. Access to check target positions.

    17:05 Vince found target ladder positions needed to be recalibrated. We were off by about 50 revolutions (~2.1 cm). Recalibrated.

    17:15 Beam back on.

    18:00 Thunderstorm causing beam problems. Lots of trips.
    Completed kinematic 11.15.1
    Now set for kinematic 11.16

    19:15 MCC reports beam off; hardware problems.
    24:00 Still no beam, latest estimate is 45 mins.
    Kevin suggests cutting remaining kinematics for polarized 3He by 40% due to time lost.